OpenStudy (anonymous):

Which function has the greatest rate of change? A. 6x – 7y = 21 B. y = 2x – 2 C. x y 2 7 3 13 4 19

OpenStudy (anonymous):

if it isnt any of those then there D but its a graph and i dont want to attach it :P im that tired

OpenStudy (compassionate):

The rate of change is the numeral change between two points. You can find the Rate of Change by finding the slope.

OpenStudy (compassionate):

For example, 

OpenStudy (igreen):

6x – 7y = 21 We can change this into slope-intercept form(y = mx + b): Subtract 6x to both sides: -7y = -6x + 21 Divide -7 to both sides: y = 6/7x + 3 Therefore, your slope for A is 6/7.

OpenStudy (igreen):

Can you tell what the slope for y = 2x - 2 is?

OpenStudy (anonymous):

2

OpenStudy (igreen):

Yep, so the slope for B is 2.

OpenStudy (igreen):

Now for 3, we can find the slope between two points from the table. Let's take (2, 7) and (3, 13): \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{13-7}{3-2}\) Subtract: \(m = \dfrac{6}{1}\) Can you divide that? @RainbowBubbles01

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Umm, the slop of y = 2x - 2 is actually -2.

OpenStudy (anonymous):

Not 2.

OpenStudy (igreen):

No it isn't.

OpenStudy (igreen):

It's 2.

OpenStudy (igreen):

Anyway, the slope of A is 6/7, B is 2, and C is 6.

OpenStudy (anonymous):

Oh, I got the slope and y-intercept mixed up. Lol.

OpenStudy (igreen):

So which one is bigger? 6/7, 2, or 6? @RainbowBubbles01

OpenStudy (anonymous):

6/7?

OpenStudy (anonymous):

thats D.

OpenStudy (igreen):

Well the slope of D is 4.. So we have 6/7, 2, 6, and 4. 6/7 is the same thing as 0.86 0.86 is the smallest one.

OpenStudy (igreen):

So which is bigger? Two, four, or six?

OpenStudy (anonymous):

6

OpenStudy (igreen):

Yep, and since C is 6, that's your answer.